# Kai Zeng    Zhanqian Wu    Kaixin Xiong    Xiaobao Wei    Xiangyu Guo    Zhenxin Zhu    Kalok Ho    Lijun Zhou    Bohan Zeng    Ming Lu    Haiyang Sun    Bing Wang    Guang Chen    Hangjun Ye    Wentao Zhang [ [ [

### v2

@@ 1 Introduction
Perception tasks such as 3D object detection (Li et al., 2022, 2024c; Wang et al., 2023a, 2025) and 3D tracking (Wang et al., 2023a; Zhang et al., 2023c; Han et al., 2025), which support planning and decision-making (Jiang et al., 2023; Hu et al., 2023), are extremely important in autonomous driving. The performance of perception models, however, is highly dependent on large-scale annotated training datasets (Caesar et al., 2020; Wang et al., 2023b). To ensure reliability in rare but critical safety scenarios, it is essential to gather adequate long-tail data. Although the autonomous driving community has developed a thorough 3D annotation pipeline to facilitate data acquisition (Zhao et al., 2025b), collecting long-tail data remains highly time-consuming and labor-intensive.
Driving world models based on diffusion model (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) have been developed in autonomous driving to generate synthetic data from scene layouts (e.g., BEV maps, 3D bounding boxes) and text (Gao et al., 2023; Wen et al., 2024; Guo et al., 2025). However, they provide limited control over the poses and appearances of individual objects, which restricts the ability to generate diverse synthetic data (Li et al., 2025). Editing methods (Singh et al., 2024; Liang et al., 2025b; Yu et al., 2025) instead of generation extend driving world models by using reference images and 3D bounding boxes to modify object appearance and pose, allowing for diverse corner cases. However, single-view insertion limits their application in multi-view BEV perception. Reconstruction-based methods using NeRF or 3DGS enable precise control over geometric structures (Chen et al., 2021; Zanjani et al., 2025). While geometry is preserved, sparse training views often lead to artifacts and incomplete renderings. Additionally, the lack of illumination modeling results in inconsistencies between inserted objects and the background.
More importantly, we argue that the data augmentation experiments of previous methods (Wen et al., 2024; Li et al., 2024a) are unfair, as they usually leverage a training strategy that first pretrains on synthetic data and finetunes on real data, resulting in twice the epochs compared to the baseline (real data only). We find that, under the same number of training epochs, large amounts of synthetic datasets offer little to no advantage and can even perform worse than using real data alone. As shown in Fig. 1, under the 2× epoch setting, models trained exclusively on real data achieve higher mAP and NDS compared to those trained on real and synthetic data. Given the importance of downstream perception tasks for autonomous driving, we believe it is essential to rethink the effectiveness of the driving world model as a synthetic data generator for these tasks.
@@
Specifically, Dream4Drive leverages a multi-view video inpainting model finetuned from the Diffusion Transformer (Peebles and Xie, 2023). Unlike prior methods that rely on sparse spatial controls (e.g., BEV maps and 3D bounding boxes), Dream4Drive uses dense 3D-aware guidance maps (Yu and Smith, 2019; Liang et al., 2025a) (e.g., depth, normal, edge, cutout, and mask) to preserve the geometry and appearance of the original video, while editing them by rendering 3D assets into these maps. This design enables instance-level, cross-view consistent video editing, ensuring both visual realism and geometric fidelity. The generated videos not only achieve superior quality but can also be directly used to train state-of-the-art perception models.
To facilitate diverse 3D-aware video editing, we design a pipeline that automatically acquires high-quality 3D assets based on a target scene image or video of a desired asset category. We first apply the image segmentation model (Ren et al., 2024; Lin et al., 2024) to localize and crop objects of the specified category, then employ the image generation model (Wu et al., 2025) to generate multi-view consistent images of the target object. These images are fed into a mesh generation model (Hunyuan3D et al., 2025) to generate high-quality 3D assets. We present DriveObj3D, a large-scale 3D asset dataset that encompasses typical categories found in driving scenarios, to support future research. Our main contributions are:
@@ 2 Related Work
Video Generation in Autonomous Driving. High-quality data is crucial for training perception models in autonomous driving, motivating growing interest in driving video generation. Early approaches (Yang et al., 2023; Gao et al., 2023; Luo et al., 2025; Li et al., 2024b) employ diffusion models with ControlNet, conditioned on BEV maps (Ma et al., 2024b) and 3D bounding boxes, to generate paired image data (An et al., 2024, 2025b; Luo et al., 2024c, a, b). Recent works (Gao et al., 2024a; Jiang et al., 2024; Li et al., 2024a; Ji et al., 2025; Guo et al., 2025) adopt powerful Diffusion Transformers (DiTs) to further enhance generation quality. SubjectDrive (Huang et al., 2024a) uses an external subject bank to enhance vehicle appearance diversity. However, these methods depend on original scene layouts, which limit geometric diversity and struggle to generate high-quality long-tail corner cases, thereby restricting their effectiveness for downstream perception.
Video Editing in Autonomous Driving. To enrich scene diversity, object-level editing methods insert new objects into existing videos using reference images and 3D bounding boxes (Singh et al., 2024; Liang et al., 2025b; Buburuzan et al., 2025; An et al., 2025a). More recent NeRF- (Mildenhall et al., 2021) and 3DGS-based (Kerbl et al., 2023) approaches (Chen et al., 2021; Huang et al., 2024b; Chen et al., 2024; Wei et al., 2024) improve geometric fidelity, yet remain limited by sparse views, inconsistent lighting, and restricted background diversity. In contrast, our approach builds on generative models and introduces a novel 3D-aware video editing mechanism, enabling seamless insertion of diverse 3D assets and producing geometrically and visually diverse data that effectively boosts downstream performance.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDMs) (Rombach et al., 2022) address the high computational cost of diffusion models by operating in a lower-dimensional latent space. Given an image $\mathbf{x}$, an encoder $\mathcal{E}$ is used to obtain the corresponding latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$. The forward diffusion process in the latent space is defined as a gradual noising process:
@@
ControlNets (Zhang et al., 2023a) extend LDMs by incorporating additional conditioning signals $\mathbf{c}$ to provide finer control over the generation process. In particular, the reverse diffusion process is modified to condition on $\mathbf{c}$:
@@
For an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we use Depth Anything (Yang et al., 2024) to obtain the depth map $D\in\mathbb{R}^{H\times W\times 1}$. The normal map $N\in\mathbb{R}^{H\times W\times 3}$ is then derived from the depth map. We also use OpenCV’s Canny edge detector to get the edge map $E\in\mathbb{R}^{H\times W\times 1}$. It is important to note that the depth, normal, and edge information within the foreground object regions are masked since our model needs to learn to generate an edited video based on the target 3D asset.
@@
The video generation process in recent methods (Wen et al., 2024; Li et al., 2024a; Guo et al., 2025) relies on sparse spatial controls, such as bird’s eye view (BEV) maps and projected 3D bounding boxes. While this approach allows for some control over the synthesized content, it often falls short in achieving high-quality generation. Instead of relying on sparse spatial controls, we fine-tune a driving world model to generate edited videos from the dense 3D-aware guidance maps, including depth map ($D$), normal map ($N$), and edge map ($E$) for the background and the image ($O$) and mask ($M$) for the foreground object.
@@
We adopt rectified flow (Liu et al., 2022) for stable sampling and classifier-free guidance (Ho and Salimans, 2022) to balance text with multiple 3D geometric conditions, thereby improving controllability. The main training objective is a simplified diffusion loss that predicts the noise component at each timestep:
@@
where $\epsilon$ is Gaussian noise, $\epsilon_{\theta}$ the predicted noise, $\mathbf{z}_{t}$ the noisy latent at timestep $t$, and $\mathbf{c}$ all conditioning signals. To achieve precise instance-level control, we introduce a Foreground Mask Loss as Ji et al. (2025) and an LPIPS loss (Zhang et al., 2018). The final training objective is:
@@ 4 DriveObj3D
Concretely, we first localize and segment the target object. Given an input image $I$ and category label class, Grounded-SAM (Ren et al., 2024) is applied to segment the object $I_{\text{target}}$. To overcome the occlusion of target object, we employ a multi-view image generation model Qwen-Image (Wu et al., 2025). Conditioned on $I_{\text{target}}$, Qwen-Image synthesizes a set of novel views $\{I_{v}\}_{v=1}^{N}$, which are subsequently fed into a multi-view reconstruction model Hunyuan3D (Hunyuan3D et al., 2025) to recover the final 3D mesh.
@@
As shown in Fig. 6, assets generated by Text-to-3D methods (Xiang et al., 2025) often exhibit style inconsistencies with the original data, while single-view approaches (Hunyuan3D et al., 2025) tend to produce incomplete assets. In contrast, our simple yet effective pipeline leverages multi-view synthesis to generate complete and high-fidelity assets, even under severe occlusions. To support large-scale downstream driving tasks, we construct a diverse asset dataset, DriveObj3D, covering a wide range of categories in driving scenarios in Fig. 7. All assets will be released publicly.
@@ 5.1 Setup
We utilize the nuScenes dataset (Caesar et al., 2020) for building 3D assets and finetuning our video generation model. It comprises 1000 scenes in total, with 700 designated for training, 150 for validation, and 150 for testing. Each scene contains a 20-second multi-view video captured by 6 cameras. More details on inserted scene and asset selection are given in the Appx. A.
@@
Following Panacea (Wen et al., 2024) and SubjectDrive (Huang et al., 2024a), we primarily evaluate how the generated data improves perceptual model performance on detection and tracking tasks. Detection metrics include nuScenes Detection Score (NDS), mean Average Precision (mAP), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Tracking metrics include Average MultiObject Tracking Accuracy (AMOTA), Precision (AMOTP), Recall, and Accuracy (MOTA).
@@ 5.2 Main Results
(28130) (+420, <2%) (28130) (Wang et al., 2024a) (Lu et al., 2024) (Gao et al., 2023) (Wen et al., 2024) (Huang et al., 2024a) (Ours)
@@
(28130) (+420, <2%) (28130) (Wen et al., 2024) (Huang et al., 2024a) (Ours)
@@
As generative models continue to advance, the ability to synthesize high-resolution videos has become achievable (Gao et al., 2024b, a). To investigate the effect of high-resolution synthetic data on downstream perception models, we further conduct experiments for detection and tracking tasks at a resolution of $512\times 768$, as reported in Tab. 3 and 4.
@@ 5.3 Ablation Studies
We observe that the source of inserted assets affects the quality of synthetic data, consistent with (Ljungbergh et al., 2025). We therefore investigate how different asset sources influence downstream perception performance.
@@
As shown in Tab. 5, although Trellis (Xiang et al., 2025) can generate high-quality assets, its style does not fully match autonomous driving scenarios, which can lead to artifacts and degraded quality when assets are inserted, negatively affecting downstream tasks. Hunyuan3D (Hunyuan3D et al., 2025)’s single-view generation also underperforms our multiview approach, since single-image assets may be incomplete, whereas our method produces complete, high-quality 3D assets.
@@ A.1 Model and Training Details
DriveObj3D Generation Pipeline. We adopt Grounded-SAMRen et al. (2024) for image segmentation, Qwen-Image-EditWu et al. (2025) for image generation, and Hunyuan3D 2.0Hunyuan3D et al. (2025) for 3D asset generation.
Inpainting Diffusion Model. For video synthesis, Dream4Drive is constructed upon a DiT-based architecture. The backbone is initialized with pretrained weights from MagicDriveDiT (Gao et al., 2024a), which originally conditioned on BEV maps and 3D bounding boxes. We extend its ControlNet branch by incorporating novel 3D-aware guidance maps, and fine-tune the model on the nuScenes dataset. The generated videos have a resolution of $512\times 768$ with 33 frames.
@@
LPIPS Loss. To further improve perceptual quality, we adopt the LPIPS (Learned Perceptual Image Patch Similarity) loss (Zhang et al., 2018), which measures feature-level perceptual differences between prediction and ground truth:
@@ Appendix C Additional Experiments on Generation Quality
DriveDream-2 Zhao et al. (2025a) 12Hz $\times512$ 55.70 11.20
@@
Controllability. The controllability of our approach is quantitatively evaluated using perception metrics from StreamPETR Wang et al. (2023a). We first generate the entire nuScenes validation set with Dream4Drive, and then measure perception performance using a pre-trained StreamPETR model. The relative metrics, compared to those obtained on real data, indicate how well the generated samples align with the original annotations. As shown in Tab. 9, Dream4Drive achieves a relative performance of 82%, demonstrating precise control over foreground object locations.
Generation Quality. As shown in Tab. 10, our method achieves FVD 31.84 and FID 5.80, outperforming layout- and 3D semantics-guided methods Gao et al. (2023, 2024a); Li et al. (2024a); Ji et al. (2025). The results indicate improved motion consistency and preserved visual fidelity, confirming that our 3D-aware guidance maps effectively generate both foreground and background content. Notably, no post-processing or selective filtering was applied to the generated videos.

### v3

@@ 1 Introduction
Perception tasks such as 3D object detection (li2022coda; li2024bevformer; wang2023exploring; wang2025survey) and 3D tracking (wang2023exploring; zhang20233d; han2025novel), which support planning and decision-making (jiang2023vad; hu2023planning), are extremely important in autonomous driving. The performance of perception models, however, is highly dependent on large-scale annotated training datasets (caesar2020nuscenes; wang2023we_asap). To ensure reliability in rare but critical safety scenarios, it is essential to gather adequate long-tail data. Although the autonomous driving community has developed a thorough 3D annotation pipeline to facilitate data acquisition (zhao2025survey), collecting long-tail data remains highly time-consuming and labor-intensive.
Driving world models based on diffusion model (LDM) and ControlNet (zhang2023adding) have been developed in autonomous driving to generate synthetic data from scene layouts (e.g., BEV maps, 3D bounding boxes) and text (gao2023magicdrive; wen2024panacea; guo2025genesis). However, they provide limited control over the poses and appearances of individual objects, which restricts the ability to generate diverse synthetic data (li2025realistic). Editing methods (singh2024genmm; liang2025driveeditor; yu2025vehiclesim) instead of generation extend driving world models by using reference images and 3D bounding boxes to modify object appearance and pose, allowing for diverse corner cases. However, single-view insertion limits their application in multi-view BEV perception. Reconstruction-based methods using NeRF or 3DGS enable precise control over geometric structures (chen2021geosim; zanjani2025gaussian). While geometry is preserved, sparse training views often lead to artifacts and incomplete renderings. Additionally, the lack of illumination modeling results in inconsistencies between inserted objects and the background.
More importantly, we argue that the data augmentation experiments of previous methods (wen2024panacea; li2024uniscene) are unfair, as they usually leverage a training strategy that first pretrains on synthetic data and finetunes on real data, resulting in twice the epochs compared to the baseline (real data only). We find that, under the same number of training epochs, large amounts of synthetic datasets offer little to no advantage and can even perform worse than using real data alone. As shown in Fig. 1, under the 2× epoch setting, models trained exclusively on real data achieve higher mAP and NDS compared to those trained on real and synthetic data. Given the importance of downstream perception tasks for autonomous driving, we believe it is essential to rethink the effectiveness of the driving world model as a synthetic data generator for these tasks.
@@
Specifically, Dream4Drive leverages a multi-view video inpainting model finetuned from the Diffusion Transformer (dit). Unlike prior methods that rely on sparse spatial controls (e.g., BEV maps and 3D bounding boxes), Dream4Drive uses dense 3D-aware guidance maps (yu2019inverserendernet; liang2025diffusion) (e.g., depth, normal, edge, cutout, and mask) to preserve the geometry and appearance of the original video, while editing them by rendering 3D assets into these maps. This design enables instance-level, cross-view consistent video editing, ensuring both visual realism and geometric fidelity. The generated videos not only achieve superior quality but can also be directly used to train state-of-the-art perception models.
To facilitate diverse 3D-aware video editing, we design a pipeline that automatically acquires high-quality 3D assets based on a target scene image or video of a desired asset category. We first apply the image segmentation model (ren2024grounded; lin2024draw) to localize and crop objects of the specified category, then employ the image generation model (wu2025qwenimagetechnicalreport) to generate multi-view consistent images of the target object. These images are fed into a mesh generation model (hunyuan3d2025hunyuan3d) to generate high-quality 3D assets. We present DriveObj3D, a large-scale 3D asset dataset that encompasses typical categories found in driving scenarios, to support future research. Our main contributions are:
@@ 2 Related Work
Video Generation in Autonomous Driving. High-quality data is crucial for training perception models in autonomous driving, motivating growing interest in driving video generation. Early approaches (yang2023bevcontrol; gao2023magicdrive; luo2025drvhierarchicalperceptiontemporalcognitionframework; li2024survey) employ diffusion models with ControlNet, conditioned on BEV maps (ma2024vision) and 3D bounding boxes, to generate paired image data (an2024mc; an2025concept; luo2024llm; luo2024panosent; luo2024nus). Recent works (gao2024magicdrivedit; jiang2024dive; li2024uniscene; ji2025cogen; guo2025genesis) adopt powerful Diffusion Transformers (DiTs) to further enhance generation quality. SubjectDrive (huang2024subjectdrive) uses an external subject bank to enhance vehicle appearance diversity. However, these methods depend on original scene layouts, which limit geometric diversity and struggle to generate high-quality long-tail corner cases, thereby restricting their effectiveness for downstream perception.
Video Editing in Autonomous Driving. To enrich scene diversity, object-level editing methods insert new objects into existing videos using reference images and 3D bounding boxes (singh2024genmm; liang2025driveeditor; buburuzan2025mobi; an2025unictokens). More recent NeRF- (mildenhall2021nerf) and 3DGS-based (kerbl20233dgs) approaches (chen2021geosim; huang2024s3g; chen2024omnire; wei2024emd) improve geometric fidelity, yet remain limited by sparse views, inconsistent lighting, and restricted background diversity. In contrast, our approach builds on generative models and introduces a novel 3D-aware video editing mechanism, enabling seamless insertion of diverse 3D assets and producing geometrically and visually diverse data that effectively boosts downstream performance.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDMs) (LDM) address the high computational cost of diffusion models by operating in a lower-dimensional latent space. Given an image $\mathbf{x}$, an encoder $\mathcal{E}$ is used to obtain the corresponding latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$. The forward diffusion process in the latent space is defined as a gradual noising process:
@@
ControlNets (controlnet2023) extend LDMs by incorporating additional conditioning signals $\mathbf{c}$ to provide finer control over the generation process. In particular, the reverse diffusion process is modified to condition on $\mathbf{c}$:
@@
For an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we use Depth Anything (yang2024depth) to obtain the depth map $D\in\mathbb{R}^{H\times W\times 1}$. The normal map $N\in\mathbb{R}^{H\times W\times 3}$ is then derived from the depth map. We also use OpenCV’s Canny edge detector to get the edge map $E\in\mathbb{R}^{H\times W\times 1}$. It is important to note that the depth, normal, and edge information within the foreground object regions are masked since our model needs to learn to generate an edited video based on the target 3D asset.
@@
The video generation process in recent methods (wen2024panacea; li2024uniscene; guo2025genesis) relies on sparse spatial controls, such as bird’s eye view (BEV) maps and projected 3D bounding boxes. While this approach allows for some control over the synthesized content, it often falls short in achieving high-quality generation. Instead of relying on sparse spatial controls, we fine-tune a driving world model to generate edited videos from the dense 3D-aware guidance maps, including depth map ($D$), normal map ($N$), and edge map ($E$) for the background and the image ($O$) and mask ($M$) for the foreground object.
@@
We adopt rectified flow (liu2022flow) for stable sampling and classifier-free guidance (ho2022classifier) to balance text with multiple 3D geometric conditions, thereby improving controllability. The main training objective is a simplified diffusion loss that predicts the noise component at each timestep:
@@
where $\epsilon$ is Gaussian noise, $\epsilon_{\theta}$ the predicted noise, $\mathbf{z}_{t}$ the noisy latent at timestep $t$, and $\mathbf{c}$ all conditioning signals. To achieve precise instance-level control, we introduce a Foreground Mask Loss as ji2025cogen and an LPIPS loss (zhang2018unreasonable). The final training objective is:
@@ 4 DriveObj3D
Concretely, we first localize and segment the target object. Given an input image $I$ and category label class, Grounded-SAM (ren2024grounded) is applied to segment the object $I_{\text{target}}$. To overcome the occlusion of target object, we employ a multi-view image generation model Qwen-Image (wu2025qwenimagetechnicalreport). Conditioned on $I_{\text{target}}$, Qwen-Image synthesizes a set of novel views $\{I_{v}\}_{v=1}^{N}$, which are subsequently fed into a multi-view reconstruction model Hunyuan3D (hunyuan3d2025hunyuan3d) to recover the final 3D mesh.
@@
As shown in Fig. 6, assets generated by Text-to-3D methods (xiang2025structured) often exhibit style inconsistencies with the original data, while single-view approaches (hunyuan3d2025hunyuan3d) tend to produce incomplete assets. In contrast, our simple yet effective pipeline leverages multi-view synthesis to generate complete and high-fidelity assets, even under severe occlusions. To support large-scale downstream driving tasks, we construct a diverse asset dataset, DriveObj3D, covering a wide range of categories in driving scenarios in Fig. 7. All assets will be released publicly.
@@ 5.1 Setup
We utilize the nuScenes dataset (caesar2020nuscenes) for building 3D assets and finetuning our video generation model. It comprises 1000 scenes in total, with 700 designated for training, 150 for validation, and 150 for testing. Each scene contains a 20-second multi-view video captured by 6 cameras. More details on inserted scene and asset selection are given in the Appx. A.
@@
Following Panacea (wen2024panacea) and SubjectDrive (huang2024subjectdrive), we primarily evaluate how the generated data improves perceptual model performance on detection and tracking tasks. Detection metrics include nuScenes Detection Score (NDS), mean Average Precision (mAP), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Tracking metrics include Average MultiObject Tracking Accuracy (AMOTA), Precision (AMOTP), Recall, and Accuracy (MOTA).
@@ 5.2 Main Results
(28130) (+420, <2%) (28130) (wang2024drivedreamer) (lu2024wovogen) (gao2023magicdrive) (wen2024panacea) (huang2024subjectdrive) (Ours)
@@
(28130) (+420, <2%) (28130) (wen2024panacea) (huang2024subjectdrive) (Ours)
@@
As generative models continue to advance, the ability to synthesize high-resolution videos has become achievable (gao2024vista; gao2024magicdrivedit). To investigate the effect of high-resolution synthetic data on downstream perception models, we further conduct experiments for detection and tracking tasks at a resolution of $512\times 768$, as reported in Tab. 3 and 4.
@@ 5.3 Ablation Studies
We observe that the source of inserted assets affects the quality of synthetic data, consistent with (ljungbergh2025r3d2). We therefore investigate how different asset sources influence downstream perception performance.
@@
As shown in Tab. 5, although Trellis (xiang2025structured) can generate high-quality assets, its style does not fully match autonomous driving scenarios, which can lead to artifacts and degraded quality when assets are inserted, negatively affecting downstream tasks. Hunyuan3D (hunyuan3d2025hunyuan3d)’s single-view generation also underperforms our multiview approach, since single-image assets may be incomplete, whereas our method produces complete, high-quality 3D assets.
@@ A.1 Model and Training Details
DriveObj3D Generation Pipeline. We adopt Grounded-SAMren2024grounded for image segmentation, Qwen-Image-Editwu2025qwenimagetechnicalreport for image generation, and Hunyuan3D 2.0hunyuan3d2025hunyuan3d for 3D asset generation.
Inpainting Diffusion Model. For video synthesis, Dream4Drive is constructed upon a DiT-based architecture. The backbone is initialized with pretrained weights from MagicDriveDiT (gao2024magicdrivedit), which originally conditioned on BEV maps and 3D bounding boxes. We extend its ControlNet branch by incorporating novel 3D-aware guidance maps, and fine-tune the model on the nuScenes dataset. The generated videos have a resolution of $512\times 768$ with 33 frames.
@@
LPIPS Loss. To further improve perceptual quality, we adopt the LPIPS (Learned Perceptual Image Patch Similarity) loss (zhang2018unreasonable), which measures feature-level perceptual differences between prediction and ground truth:
@@ Appendix C Additional Experiments on Generation Quality
DriveDream-2 zhao2025drivedreamer 12Hz $\times512$ 55.70 11.20
@@
Controllability. The controllability of our approach is quantitatively evaluated using perception metrics from StreamPETR wang2023exploring. We first generate the entire nuScenes validation set with Dream4Drive, and then measure perception performance using a pre-trained StreamPETR model. The relative metrics, compared to those obtained on real data, indicate how well the generated samples align with the original annotations. As shown in Tab. 9, Dream4Drive achieves a relative performance of 82%, demonstrating precise control over foreground object locations.
Generation Quality. As shown in Tab. 10, our method achieves FVD 31.84 and FID 5.80, outperforming layout- and 3D semantics-guided methods gao2023magicdrive; gao2024magicdrivedit; li2024uniscene; ji2025cogen. The results indicate improved motion consistency and preserved visual fidelity, confirming that our 3D-aware guidance maps effectively generate both foreground and background content. Notably, no post-processing or selective filtering was applied to the generated videos.